# Depth-Width Tradeoffs in Algorithmic Reasoning of Graph Tasks with Transformers

### v1

@@ 6.2 Critical Width
6.3 Adjacency Rows Graph Tokenization
@@
We compared the adjacency-row representation to the edge-list representation by training a transformer model on three Open Graph Benchmark (OGB) (Hu et al., 2020) datasets: ogbg-molhiv, ogbg-molbbbp, and ogbg-molbace.
In ogbg-molhiv, the task is to predict whether a molecule inhibits HIV replication, a binary classification task based on molecular graphs with atom-level features and bond-level edge features.
ogbg-bbbp involves predicting blood-brain barrier permeability, a crucial property for drug development.
ogbg-bace focuses on predicting the ability of a molecule to bind to the BACE1 enzyme, associated with Alzheimer’s disease.
We also evaluated a tokenization using the eigenvectors and eigenvalues of the graph Laplacian, as commonly used in the literature (Dwivedi & Bresson, 2021; Kreuzer et al., 2021). More experimental details, including the dataset statistics can be found in Appendix E.
The results of our evaluation are summarized in Table 1. In all three tasks, the adj-rows representation achieved better ROC-AUC scores than the edge-list representation.
In two out of the three, it also improved upon the commonly used Laplacian eigenvectors representations.
The results suggest that the adjacency representation we use in this paper is empirically effective, and should be considered alongside the commonly used Laplacian eigenvector representation.
Table 1: This table compares the ROC-AUC performance metrics for different graph representations: Edge List, Adjacency Rows, and Laplacian Eigenvectors (LE).
The results are reported as the average ROC-AUC over three random seeds, along with the corresponding standard deviations.
@@ 7 Discussion and Future Work
This paper uses a collection of graph algorithmic tasks—including subgraph detection, connectivity, and Eulerian cycle verification—to demonstrate the powers of width bounded-depth transformers that take as input node adjacency encodings. These results include sharp theoretical thresholds that demonstrate the trade-offs between constant, linear, quadratic, and intermediate width regimes. Our empirical results validate the efficiency and accuracy of our choice of scaling regime and embedding strategy.
There are numerous possible extensions of this work.
One future direction is to study different graph tokenization schemes, beyond the node-adjacency encoding of this work and edge-list encoding of (Sanford et al., 2024a).
A particularly notable alternative is the smallest eigenvectors of the graph Laplacian, presented as a vector of components for each node.
This spectral embedding is a standard embedding scheme for GNNs, and the techniques and tasks developed in this paper would likely be relevant to proving similar bounds.
We provide a preliminary exploration of the trade-offs between the intrinsically local characteristics of adjacency-based tokenization schemes and more global spectral approaches in Appendix A.
Another future direction is to study the optimization and generalization capabilities of transformers to solve graph problems, beyond the expressiveness limitations presented in this work.
@@ Appendix A Alternative tokenization approaches
The node-adjacency tokenization is amenable for analyzing local structures around each node. Most simply, the degree of each node can be computed in a sequence-wise manner with node-adjacency tokenization with embedding dimension $m=1$ by simply computing the inner products $\langle\boldsymbol{1}_{n},\mathbf{x}_{i}\rangle$. Constructions like Theorems 4.3 and 4.4 further demonstrate the abilities of adjacency encodings to aggregate local structures.
In contrast, choosing the smallest eigenvectors in the alternative encoding makes it impossible to even compute each node degree without having embedding dimension $m$ growing linearly in the node count $n$222 Consider the task of computing the degree of a particular node of a graph consisting of $\frac{n}{3}$ disconnected linear subgraphs, each with three nodes connected by two edges. The zero eigenvalue thus has multiplicity $\frac{n}{3}$, and hence the eigenvectors $\mathbf{v}_{1},\dots,\mathbf{v}_{n/3}$ exist solely as indicators of connected components. Therefore, if nodes $i,j,k$ comprise a cluster and $m\leq\frac{n}{3}$, then their embeddings $\mathbf{y}_{i},\mathbf{y}_{j},\mathbf{y}_{k}$ are identical. .
@@
Appendix B Proofs from Section 4
B.1 Proof of Theorem 4.1
@@ Appendix B Proofs from Section 4
We now define the construction of the transformer. The input to the transformer are adjacency rows, where we concatenate positional encodings that include the row number. Namely, the $i$-th input token is equal to $\begin{pmatrix}\mathbf{x}_{i}\\ i\end{pmatrix}$, where $\mathbf{x}_{i}$ is the $i$-th row of the adjacency matrix of the graph. The first layer of self-attention will not effect the inputs. This can be thought of as choosing $V=\boldsymbol{0}$ (while $K$ and $Q$ are arbitrary), and using the residual connection so that the tokens remain the same as the input. We now use Lemma B.1 to construct a $3$-layer MLP that changes the embedding of each token such that it includes for each node its neighbors. The MLP will does not change the positional encoding, this can be done since ReLU network can simulate the identity function by $z\mapsto\sigma(z)-\sigma(-z)$. We add another layer to the MLP that maps $\mathbb{R}^{3}\ni\mathbf{v}_{i}\mapsto\mathbf{u}_{i}\in\mathbb{R}^{3n}$, where $(\mathbf{u}_{i})_{3(i-1)+1:3i}=\mathbf{v}_{i}$ and all the other entries of $\mathbf{u}_{i}$ are equal to 00.
@@
Finally, we use the MLP to determine whether the input graph, whose embedded as a list of edges, is connected. This can be done by an MLP since it has the universal approximation property (Cybenko, 1989; Leshno et al., 1993). Although we don’t specify the exact size of this MLP, it can be bounded since there are efficient deterministic algorithms for determining connectivity. These algorithms can be simulated using ReLU networks.
Note that the output of the connectivity problem is either 00 or $1$, thus it is enough to approximate a solution of this task up to a constant error (say, of $\frac{1}{4}$), and then use another layer to threshold over the answer.
@@
This is a linear combination of indicators, where each indicator can be implemented by the function $f(z)=\sigma(x)-\sigma(x-1)$. Note that the input to the MLP is either 00 or $1$ in each coordinate, thus the output of this function will be $\begin{pmatrix}i+j\\ i^{2}+j^{2}\end{pmatrix}$. We have that $i+j$ and $i^{2}+j^{2}$ determine the values of $i$ and $j$. This means that if $i_{1}+j_{1}=i_{2}+j_{2}$ and $i_{1}^{2}+j_{1}^{2}=i_{2}^{2}+j_{2}^{2}$ then $i_{1}=i_{2}$ or $i_{1}=j_{2}$ and similarly for $i_{2}$. Since there are $O(n^{2})$ different possible values, we can construct a network with $2$-layers and $O(n^{2})$ width that outputs $\begin{pmatrix}i\\ j\end{pmatrix}$ up to changing the order of $i$ and $j$.
B.2 Proof of Theorem 4.2
@@
Our proof relies on a communication complexity lower bound for the set disjointness problem, and is similar to the arguments from (Sanford et al., 2024c; Yehudai et al., 2024). The lower bound for communication complexity is the following: Alice and Bob are given inputs $a,b\in\{0,1\}^{s}$ respectively, and their goal is to find $\max a_{i}b_{i}$ by sending single bit-messages to each other in a sequence of communication rounds. The lower bound says that any deterministic protocol for solving such a task requires at least $s$ rounds of communication.
@@
By the equivalence between the set disjointness and the directed $2$-cycle that was described above, Alice returns $1$ iff the inputs $\max_{i}a_{i}b_{i}=1$, and 00 otherwise. The total number of bits transmitted in this protocol in the case there are residual connection is $O(nmpHL)$, since there are $O(nmpH)$ bits transferred in each layer. The lower bound is determined by the size of the input which is $s=n^{2}$, hence $mpHL=\Omega(n)$. In the case there are no residual connections there is no dependence on $L$, hence the lower bound becomes $mpH=\Omega(n)$. ∎
B.2.1 An extension to bounded degree graphs
In order to prove the optimality result in Theorem 4.4 for the task of determining the existence of a 2-cycle in bounded-degree graphs, we state the following theorem.
@@ B.2.1 An extension to bounded degree graphs
The proof is nearly identical to that of Theorem 4.2, except that we alter the reduction to ensure that that the graph possessed by Alice and Bob is of degree $d$.
@@
where $c>0$ is some sufficiently large constant the determines the temperature of the softmax. We first have that $X^{\top}K^{\top}QX=A$. Since all the values of $A$ are either 00 or $1$, for a sufficiently large $c>0$, the softmax behave similarly to the hardmax and we get that:
$VX\text{sm}(A)=\begin{pmatrix}A^{2}\cdot\text{deg}(A)^{-1}\\ A\cdot\text{deg}(A)^{-1}\\ \boldsymbol{0}_{d\times d}\end{pmatrix}~{},$
@@
Constructing such an MLP can be done by calculating the degree of each token from $A\cdot\text{deg}(A)^{-1}$ and multiplying the first $d$ coordinates of each token by this degree. This can be done since the entries of this matrix is either 00 or the inverse of the degree of node $i$, thus it requires only inverting an integer between $1$ and $n$. By Lemma B.3 this can be done by a $2$-layer MLP with width $n$. This finishes the construction for calculating $A^{2}$.
@@
The optimality result is proved using the same methodology as Theorem 4.2. It is stated and proved formally as Theorem B.2.
The proof of the construction adapts an argument from Theorem 2 of (Sanford et al., 2024c), which shows that a sparse averaging task can be solved with bounded-width transformers. We make use of the following fact, which is a consequence of the Restricted Isometry Property analysis of (Candes & Tao, 2005; Mendelson et al., 2005).
@@
We use this fact to prove Theorem 4.4.
@@
By applying Lemma B.4, we find that
@@
$(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{j,i}$ $\leq\frac{3}{2}c,\quad\text{otherwise.}$.
@@
Thus, for sufficiently large $c$, all nonzero elements (after rounding) of $\text{sm}(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{\cdot,i}$ belongs to indices $j\in[n]$ if there exist at least one 2-cycle containing node $i$; if not, then $\text{sm}(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{n+1,i}=1$ and all others are zero.
@@
Appendix C Proofs from Section 5
C.1 Proof of Theorem 5.1
See 5.1
@@ Appendix C Proofs from Section 5
We define $\tilde{\mathbf{x}}_{i}\in\mathbb{R}^{n^{2-2/k}}$ for $i\equiv j(\text{mod }n^{1-1/k})$ to be equal to $\bar{\mathbf{x}}_{i}$ in the coordinates $(j-1)n^{1-1/k}+1$ until $jn^{1-1/k}$ and all the other coordinates are 00. These vectors will later be summed together across all nodes in the same set, which provides an encoding of all the edges in the set. We also define $\mathbf{w}_{i}=\mathbf{e}_{j}\in\mathbb{R}^{n^{1/k}}$ for $(j-1)n^{1/k}+1\leq i\leq jn^{1/k}$ and $\mathbf{z}_{i}=\mathbf{e}_{i}\in\mathbb{R}^{n^{1/k}}$ for $i=1,\dots,n^{1/k}$ and $\mathbf{z}_{i}=\boldsymbol{0}$ otherwise.
@@
Given two vectors $\begin{pmatrix}\tilde{\mathbf{x}}_{i}\\ \mathbf{w}_{i}\\ \mathbf{z}_{i}\\ i\end{pmatrix},\begin{pmatrix}\tilde{\mathbf{x}}_{j}\\ \mathbf{w}_{j}\\ \mathbf{z}_{j}\\ j\end{pmatrix}$, which are outputs of the previous layer, we have that: $\begin{pmatrix}\tilde{\mathbf{x}}_{i}\\ \mathbf{w}_{i}\\ \mathbf{z}_{i}\\ i\end{pmatrix}^{\top}K^{\top}Q\begin{pmatrix}\tilde{\mathbf{x}}_{j}\\ \mathbf{w}_{j}\\ \mathbf{z}_{j}\\ j\end{pmatrix}=\left\langle\mathbf{z}_{i},\mathbf{w}_{j}\right\rangle$. This shows that the first $n^{1/k}$ tokens, which represent each of the $n^{1/k}$ sets, will attend with a similar weight to every node in their set. After applying the $V$ matrix we use the residual connection only for the positional embedding vectors [^3] It is always possible to use the residual connection to affect only a subset of the coordinates. This can be done by doubling the number of unaffected coordinates, using the $V$ matrix to move the unaffected entries to these new coordinates, and then using a $1$-layer MLP to move the unaffected entries to their previous coordinates (which now include what was added through the residual connection). We omit this construction from here for brevity and since it only changes the embedding dimension by a constant factor. (namely, the last $12n^{1/k}+$ 1 coordinates). Thus, the output of the self-attention layer for the first $n^{1/k}$ tokens encodes all the edges in their set in their first $n^{2-2/k}$ coordinates. This encoding is such that there is $1$ in the $i$-th coordinates if there is an edge between nodes $v_{s}$ and $v_{r}$ in the set for where $r$ and $s$ are the unique integers such that $i\equiv r(\text{mod }n^{1-1/k})$ and $(s-1)n^{2-2/k}+1<i<sn^{2-2/k}$. Thus, the output of the self-attention layer can be written as $\begin{pmatrix}\mathbf{y}_{i}\\ \mathbf{w}_{i}\\ \mathbf{z}_{i}\\ i\end{pmatrix}\in\mathbb{R}^{n^{2-1/k}+2n^{1/k}+1}$, where $\mathbf{y}_{i}$ is either an encoding as described above (for $i\leq n^{1/k}$) or some other vector (for $i\geq n^{1/k}$) for which its exact value will not matter. The vector $\mathbf{w}_{i}$ is a positional embedding that is not needed anymore and will be removed by the MLP, and $\mathbf{z}_{i}=\mathbf{e}_{i}$ for $i\leq n^{1/k}$ and $\mathbf{z}_{i}=\boldsymbol{0}$ otherwise.
We will now construct the MLP of the second layer. First, note that given $n^{1/k}$ sets, the number of all $k$ combinations of such sets is bounded by $\binom{n^{1/k}}{k}\leq n^{k\cdot 1/k}=n$. Denote all possible combinations by $B_{1},\dots,B_{n}$ and let $\mathbf{v}_{1},\dots,\mathbf{v}_{n}\in\mathbb{R}^{n^{1/k}}$ such that $(\mathbf{v}_{i})_{j}=1$ if $B_{i}$ includes the $j$-th set, and 00 otherwise. These vectors encode all the possible combinations of such sets. The MLP will apply the following map:
@@
Finally, the MLP will be used to detect whether the given subgraph of size $k$ appears as a subgraph in the input graph (which is an encoding of the edges). The output of the MLP will be $1$ if the subgraph appears and 00 otherwise.
Note that any subgraph of size $k$ must appear in one of those combination of sets. Thus, by summing all the tokens, if their sum is greater than 00 the subgraph of size $k$ appears as a subgraph of $G$.
C.2 Proof of Theorem 5.3
@@
    every edge in $E$ appears exactly two pairs in $P$; and
@@
Consider some transformer $T$ with depth $L$ and embedding dimension $m$ that solves the Eulerian cycle verification problem for any directed multi-graph with $n$ nodes and at most $n^{2}$ edges. We use this to construct a transformer $\mathfrak{T}$ with depth $L+O(1)$ and embedding dimension $O(m+n^{1.1})$ that distinguishes between an undirected444 The one-cycle vs two-cycle conjecture applies only to undirected graphs. cycle graph of size $N$ and two cycles of size $\frac{N}{2}$ for $N=\frac{n^{2}}{2}$. The claim of the theorem follows as an immediate consequence of the one-cycle vs two-cycle conjecture (as stated in Conjecture 13 of Sanford et al. (2024a)).
@@
Therefore, if $\mathfrak{G}$ is contains a single cycle of length $N$, then $P$ verifies an Eulerian cycle in $G$ that includes all $n^{2}$ edges Otherwise, if $\mathfrak{G}$ has two cycles of length $2N$, then $P$ represents three cyclic paths, one of length $\frac{n^{2}}{2}$ and two of length $\frac{n^{2}}{4}$. Hence, there is a one-to-one correspondence between the one-cycle vs two-cycle detection problem on $\mathfrak{G}$ and the Eulerian cycle verification problem on $G$ and $P$.

### v2

@@ 6.2 Minimum width for substructure counting
In this subsection, we show that the transformer width required to fit the substructure counting tasks increases sub-quadratically with the number of nodes, as argued in  Section˜5.2. The experiments considered graphs with increasing numbers of nodes, ranging from 50 to 400 in increments of 50, and transformer widths varying from 100 to 800 in increments of 100. For each combination of graph size and transformer width, we determined the critical width at which the model failed to fit the data. The critical width is defined as the width where the training loss plateaued at more than 0.05. To determine the critical width, we conducted a grid search over each combination of graph size and model width, and selected the model that fitted the data best. The hyper-parameters we considered are provided in Appendix˜E.
To isolate the effect of width, we used a transformer model with one layer. We used two attention heads to ensure there exists a width for which the model can fit the data for all the evaluated graph sizes. The results for the 4-Cycle count task are presented in Figure˜3. Due to space limitations, the results for the triangle count task are deferred to Appendix˜D. Figure˜3 shows that the critical width increases roughly linearly with the graph size.
6.3 Comparison of graph encodings
@@
We compared the adjacency-row representation to the edge-list representation by training a transformer model on three Open Graph Benchmark (OGB) Hu et al. (2020) datasets: ogbg-molhiv, ogbg-molbbbp, and ogbg-molbace. In ogbg-molhiv, the task is to predict whether a molecule inhibits HIV replication, a binary classification task based on molecular graphs with atom-level features and bond-level edge features. ogbg-bbbp involves predicting blood-brain barrier permeability, a crucial property for drug development. ogbg-bace focuses on predicting the ability of a molecule to bind to the BACE1 enzyme, associated with Alzheimer’s disease. We also evaluated a tokenization using the eigenvectors and eigenvalues of the graph Laplacian, as commonly used in the literature (Dwivedi and Bresson, 2021; Kreuzer et al., 2021). More experimental details, including the dataset statistics can be found in Appendix˜E.
Table 1: ROC-AUC performance metrics for different graph representations: Edge List, Adjacency Rows, and Laplacian Eigenvectors (LE), averaged over 3 random seeds.
@@
The results of our evaluation are summarized in Table 1. In all three tasks, the adj-rows representation achieved better ROC-AUC scores than the edge-list representation. In two out of the three, it also improved upon the commonly used Laplacian eigenvectors representations. The results suggest that the adjacency representation we use in this paper is empirically effective, and should be considered alongside the commonly used Laplacian eigenvector representation.
@@ 7 Discussion and Future Work
This paper uses a collection of graph algorithmic tasks—including subgraph detection, one vs two cycles, and Eulerian cycle verification—to demonstrate the powers of width bounded-depth transformers that take as input node adjacency encodings. These results include sharp theoretical thresholds that demonstrate the trade-offs between constant, linear, quadratic, and intermediate width regimes. Our empirical results validate the efficiency and accuracy of our choice of scaling regime and embedding strategy.
There are numerous possible extensions of this work. One future direction is to study different graph tokenization schemes, beyond the node-adjacency encoding of this work and edge-list encoding of Sanford et al. (2024a). A particularly notable alternative is the smallest eigenvectors of the graph Laplacian, presented as a vector of components for each node. This spectral embedding is a standard embedding scheme for GNNs, and the techniques and tasks developed in this paper would likely be relevant to proving similar bounds. We provide a preliminary exploration of the trade-offs between the intrinsically local characteristics of adjacency-based tokenization schemes and more global spectral approaches in Appendix˜A. Another future direction is to study the optimization and generalization capabilities of transformers to solve graph problems, beyond the expressiveness results presented in this work.
@@ Appendix A Alternative tokenization approaches
The node-adjacency tokenization is amenable for analyzing local structures around each node. Most simply, the degree of each node can be computed in a sequence-wise manner with node-adjacency tokenization with embedding dimension $m=1$ by simply computing the inner products $\langle\boldsymbol{1}_{n},\mathbf{x}_{i}\rangle$. Constructions like Theorems˜4.3 and 4.4 further demonstrate the abilities of adjacency encodings to aggregate local structures.
In contrast, choosing the smallest eigenvectors in the alternative encoding makes it impossible to even compute each node degree without having embedding dimension $m$ growing linearly in the node count nn [^5] Consider the task of computing the degree of a particular node of a graph consisting of $\frac{n}{3}$ disconnected linear subgraphs, each with three nodes connected by two edges. The zero eigenvalue thus has multiplicity $\frac{n}{3}$, and hence the eigenvectors $\mathbf{v}_{1},\dots,\mathbf{v}_{n/3}$ exist solely as indicators of connected components. Therefore, if nodes $i,j,k$ comprise a cluster and $m\leq\frac{n}{3}$, then their embeddings $\mathbf{y}_{i},\mathbf{y}_{j},\mathbf{y}_{k}$ are identical. .
@@
Appendix B Proofs from section˜4
B.1 The $1$ vs. $2$ cycle conjecture
The most notable conjecture in distributed computing is the $1$ vs. $2$ cycle conjecture, which is a common method for providing conditional lower bounds in the MPC model. See Definition 2 in Sanford et al. [2024a] for a $(\gamma,\delta)$-MPC protocol. In simple words, this is a distributed computing protocol, where the input has length $n$, which is distributed to $\Theta(n^{1+\gamma-\delta})$ machines, each having memory of $s:=\Theta(n^{\delta})$. Each communication round, each machine calculates an arbitrary local message bounded by size $s$, and all the messages are sent simultaneously. The $1$ vs. $2$ cycle conjecture is the following:
Conjecture B.1 (Conjecture 13 in Sanford et al. [2024a]).
For any $\gamma>0$ and $\delta\in(0,1)$, any $(\gamma,\delta)$-MPC protocol that distinguishes a single length-$n$ cycle from two disjoint length-$n/2$ uses $\Omega(n)$ communication rounds.
B.2 Proof of Theorem˜4.1
@@ Appendix B Proofs from section˜4
We now define the construction of the transformer. The input to the transformer are adjacency rows, where we concatenate positional encodings that include the row number. Namely, the $i$-th input token is equal to $\begin{pmatrix}\mathbf{x}_{i}\\ i\end{pmatrix}$, where $\mathbf{x}_{i}$ is the $i$-th row of the adjacency matrix of the graph. The first layer of self-attention will not effect the inputs. This can be thought of as choosing $V=\boldsymbol{0}$ (while $K$ and $Q$ are arbitrary), and using the residual connection so that the tokens remain the same as the input. We now use Lemma˜B.2 to construct a $3$-layer MLP that changes the embedding of each token such that it includes for each node its neighbors. The MLP does not change the positional encoding, this can be done since ReLU networks can simulate the identity function by $z\mapsto\sigma(z)-\sigma(-z)$, where $\sigma$ represents the ReLU function. We add another layer to the MLP that maps $\mathbb{R}^{3}\ni\mathbf{v}_{i}\mapsto\mathbf{u}_{i}\in\mathbb{R}^{3n}$, where $(\mathbf{u}_{i})_{3(i-1)+1:3i}=\mathbf{v}_{i}$ and all the other entries of $\mathbf{u}_{i}$ are equal to 0.
@@
Finally, we use the MLP to determine whether the input graph, whose edges are embedded as a list of edges, is connected. This can be done by an MLP since it has the universal approximation property Cybenko [1989], Leshno et al. [1993]. Although we don’t specify the exact size of this MLP, it can be bounded since there are efficient deterministic algorithms for determining connectivity. These algorithms can be simulated using ReLU networks.
Note that the output of the connectivity problem is either 0 or $1$, thus it is enough to approximate a solution of this task up to a constant error (say, of $\frac{1}{4}$), and then use another layer to threshold over the answer.
@@
This is a linear combination of indicators, where each indicator can be implemented by the function $f(z)=\sigma(x)-\sigma(x-1)$. Note that the input to the MLP is either 0 or $1$ in each coordinate, thus the output of this function will be $\begin{pmatrix}i+j\\ i^{2}+j^{2}\end{pmatrix}$. We have that $i+j$ and $i^{2}+j^{2}$ determine the values of $i$ and $j$. This means that if $i_{1}+j_{1}=i_{2}+j_{2}$ and $i_{1}^{2}+j_{1}^{2}=i_{2}^{2}+j_{2}^{2}$ then $i_{1}=i_{2}$ or $i_{1}=j_{2}$ and similarly for $i_{2}$. Since there are $O(n^{2})$ different possible values, we can construct a network with $2$-layers and $O(n^{2})$ width that outputs $\begin{pmatrix}i\\ j\end{pmatrix}$ up to changing the order of $i$ and $j$.
B.3 Proof of Theorem˜4.2
@@
Our proof relies on a communication complexity lower bound for the set disjointness problem, and is similar to the arguments from Sanford et al. [2024c], Yehudai et al. [2024]. The lower bound for communication complexity is the following: Alice and Bob are given inputs $a,b\in\{0,1\}^{s}$ respectively, and their goal is to find $\max a_{i}b_{i}$ by sending single bit-messages to each other in a sequence of communication rounds. The lower bound says that any deterministic protocol for solving such a task requires at least $s$ rounds of communication.
@@
By the equivalence between the set disjointness and the directed $2$-cycle that was described above, Alice returns $1$ iff the inputs $\max_{i}a_{i}b_{i}=1$, and 0 otherwise. The total number of bits transmitted in this protocol in the case there are residual connection is $O(nmpHL)$, since there are $O(nmpH)$ bits transferred in each layer. The lower bound is determined by the size of the input which is $s=n^{2}$, hence $mpHL=\Omega(n)$. In the case there are no residual connections there is no dependence on $L$, hence the lower bound becomes $mpH=\Omega(n)$. ∎
B.3.1 An extension to bounded degree graphs
In order to prove the optimality result in Theorem˜4.4 for the task of determining the existence of a 2-cycle in bounded-degree graphs, we state the following theorem.
@@ B.3.1 An extension to bounded degree graphs
The proof is nearly identical to that of Theorem˜4.2, except that we alter the reduction to ensure that that the graph possessed by Alice and Bob is of degree $d$.
@@
where $c>0$ is some sufficiently large constant the determines the temperature of the softmax. We first have that $X^{\top}K^{\top}QX=A$. Since all the values of $A$ are either 0 or $1$, for a sufficiently large $c>0$, the softmax behave similarly to the hardmax and we get that:
$VX\textsc{sm}(A)=\begin{pmatrix}A^{2}\cdot\text{deg}(A)^{-1}\\ A\cdot\text{deg}(A)^{-1}\\ \boldsymbol{0}_{d\times d}\end{pmatrix}~$,
@@
Constructing such an MLP can be done by calculating the degree of each token from $A\cdot\text{deg}(A)^{-1}$ and multiplying the first $d$ coordinates of each token by this degree. This can be done since the entries of this matrix is either 0 or the inverse of the degree of node $i$, thus it requires only inverting an integer between $1$ and $n$. By Lemma˜B.4 this can be done by a $2$-layer MLP with width $n$. This finishes the construction for calculating $A^{2}$.
@@
The optimality result is proved using the same methodology as Theorem˜4.2. It is stated and proved formally as Theorem˜B.3.
The proof of the construction adapts an argument from Theorem 2 of Sanford et al. [2024c], which shows that a sparse averaging task can be solved with bounded-width transformers. We make use of the following fact, which is a consequence of the Restricted Isometry Property analysis of Candes and Tao [2005], Mendelson et al. [2005].
@@
We use this fact to prove Theorem˜4.4.
@@
By applying Lemma˜B.5, we find that
@@
$(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{j,i}$ $\leq\frac{3}{2}c,\quad\text{otherwise.}$
@@
Thus, for sufficiently large $c$, all nonzero elements (after rounding) of $\textsc{sm}(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{\cdot,i}$ belongs to indices $j\in[n]$ if there exist at least one 2-cycle containing node $i$; if not, then $\textsc{sm}(\tilde{X}^{\top}K^{\top}Q\tilde{X})_{n+1,i}=1$ and all others are zero.
@@
Appendix C Proofs from Section˜5
C.1 Proof of Theorem˜5.2
See 5.2
@@ Appendix C Proofs from Section˜5
We define $\tilde{\mathbf{x}}_{i}\in\mathbb{R}^{n^{2-2/k}}$ for $i\equiv j(\text{mod }n^{1-1/k})$ to be equal to $\bar{\mathbf{x}}_{i}$ in the coordinates $(j-1)n^{1-1/k}+1$ until $jn^{1-1/k}$ and all the other coordinates are 0. These vectors will later be summed together across all nodes in the same set, which provides an encoding of all the edges in the set. We also define $\mathbf{w}_{i}=\mathbf{e}_{j}\in\mathbb{R}^{n^{1/k}}$ for $(j-1)n^{1/k}+1\leq i\leq jn^{1/k}$ and $\mathbf{z}_{i}=\mathbf{e}_{i}\in\mathbb{R}^{n^{1/k}}$ for $i=1,\dots,n^{1/k}$ and $\mathbf{z}_{i}=\boldsymbol{0}$ otherwise.
@@
Given two vectors $\begin{pmatrix}\tilde{\mathbf{x}}_{i}\\ \mathbf{w}_{i}\\ \mathbf{z}_{i}\\ i\end{pmatrix},\begin{pmatrix}\tilde{\mathbf{x}}_{j}\\ \mathbf{w}_{j}\\ \mathbf{z}_{j}\\ j\end{pmatrix}$, which are outputs of the previous layer, we have that: $\begin{pmatrix}\tilde{\mathbf{x}}_{i}\\ \mathbf{w}_{i}\\ \mathbf{z}_{i}\\ i\end{pmatrix}^{\top}K^{\top}Q\begin{pmatrix}\tilde{\mathbf{x}}_{j}\\ \mathbf{w}_{j}\\ \mathbf{z}_{j}\\ j\end{pmatrix}=\left\langle\mathbf{z}_{i},\mathbf{w}_{j}\right\rangle$. This shows that the first $n^{1/k}$ tokens, which represent each of the $n^{1/k}$ sets, will attend with a similar weight to every node in their set. After applying the $V$ matrix we use the residual connection only for the positional embedding vectors [^6] It is always possible to use the residual connection to affect only a subset of the coordinates. This can be done by doubling the number of unaffected coordinates, using the $V$ matrix to move the unaffected entries to these new coordinates, and then using a $1$-layer MLP to move the unaffected entries to their previous coordinates (which now include what was added through the residual connection). We omit this construction from here for brevity and since it only changes the embedding dimension by a constant factor. (namely, the last $2n^{1/k}+1$ coordinates). Thus, the output of the self-attention layer for the first $n^{1/k}$ tokens encodes all the edges in their set in their first $n^{2-2/k}$ coordinates. This encoding is such that there is $1$ in the $i$-th coordinates if there is an edge between nodes $v_{s}$ and $v_{r}$ in the set for where $r$ and $s$ are the unique integers such that $i\equiv r(\text{mod }n^{1-1/k})$ and $(s-1)n^{2-2/k}+1<i<sn^{2-2/k}$. Thus, the output of the self-attention layer can be written as $\begin{pmatrix}\mathbf{y}_{i}\\ \mathbf{w}_{i}\\ \mathbf{z}_{i}\\ i\end{pmatrix}\in\mathbb{R}^{n^{2-1/k}+2n^{1/k}+1}$, where $\mathbf{y}_{i}$ is either an encoding as described above (for $i\leq n^{1/k}$) or some other vector (for $i\geq n^{1/k}$) for which its exact value will not matter. The vector $\mathbf{w}_{i}$ is a positional embedding that is not needed anymore and will be removed by the MLP, and $\mathbf{z}_{i}=\mathbf{e}_{i}$ for $i\leq n^{1/k}$ and $\mathbf{z}_{i}=\boldsymbol{0}$ otherwise.
We will now construct the MLP of the second layer. First, note that given $n^{1/k}$ sets, the number of all $k$ combinations of such sets is bounded by $\binom{n^{1/k}}{k}\leq n^{k\cdot 1/k}=n$. Denote all possible combinations by $B_{1},\dots,B_{n}$ and let $\mathbf{v}_{1},\dots,\mathbf{v}_{n}\in\mathbb{R}^{n^{1/k}}$ such that $(\mathbf{v}_{i})_{j}=1$ if $B_{i}$ includes the $j$-th set, and 0 otherwise. These vectors encode all the possible combinations of such sets. The MLP will apply the following map:
@@
Finally, the MLP will be used to detect whether the given subgraph of size $k$ appears as a subgraph in the input graph (which is an encoding of the edges). The output of the MLP will be $1$ if the subgraph appears and 0 otherwise.
Note that any subgraph of size $k$ must appear in one of those combination of sets. Thus, by summing all the tokens, if their sum is greater than 0 the subgraph of size $k$ appears as a subgraph of $G$.
C.2 Proof of Theorem˜5.1
@@
    every edge in $E$ appears in exactly two pairs in $P$; and
@@
Consider some transformer $T$ with depth $L$ and embedding dimension $m$ that solves the Eulerian cycle verification problem for any directed multi-graph with $n$ nodes and at most $n^{2}$ edges. We use this to construct a transformer $\mathfrak{T}$ with depth $L+O(1)$ and embedding dimension $O(m+n^{1.1})$ that distinguishes between an undirected [^7] The $1$ vs. $2$ cycle conjecture applies only to undirected graphs. cycle graph of size $N$ and two cycles of size $\frac{N}{2}$ for $N=\frac{n^{2}}{2}$. The claim of the theorem follows as an immediate consequence of the $1$ vs. $2$ cycle conjecture (as stated in Conjecture 13 of Sanford et al. [2024a]).
@@
Therefore, if $\mathfrak{G}$ is contains a single cycle of length $N$, then $P$ verifies an Eulerian cycle in $G$ that includes all $n^{2}$ edges Otherwise, if $\mathfrak{G}$ has two cycles of length $2N$, then $P$ represents three cyclic paths, one of length $\frac{n^{2}}{2}$ and two of length $\frac{n^{2}}{4}$. Hence, there is a one-to-one correspondence between the $1$ vs. $2$ cycle detection problem on $\mathfrak{G}$ and the Eulerian cycle verification problem on $G$ and $P$.